OpenStudy (anonymous):

what percentage of 30 = 12

OpenStudy (whpalmer4):

A percentage is just a ratio with 100. \(x\% = x/100\) That means our problem is simply \[\frac{12}{30} = \frac{x}{100}\]You can solve that by cross-multiplication: \[12*100 = 30*x\]Now divide both sides by 30 to get the value of \(x\).

OpenStudy (anonymous):

I was just going to say that. you beat me to the punch. Good Job.

OpenStudy (anonymous):

40%